OpenStudy (mouth4war):

What is the expression in radical form? (3p^3q) 3/4

OpenStudy (mathmale):

What part of this problem are you able to do yourself with confidence, and what part do you need help with?

OpenStudy (mouth4war):

all of it idk i dont understand it

OpenStudy (mathmale):

Explain in your own words what you're supposed to do here. Note the following:\[a ^{\frac{ b }{ c }}=\sqrt[c]{a^b}\]

OpenStudy (mathmale):

Which one involves a "fractional exponent?" Which one involves a "radical?"

OpenStudy (mathmale):

If you're unsure, why not look up "fractional exponent" and "radical" on the 'Net? You'll be seeing these terms again and again in Algebra.

OpenStudy (mouth4war):

radical means the "v" sign right?

OpenStudy (mathmale):

I get y our drift. Properly written, a radical operator looks like

OpenStudy (mathmale):

so you need to convert (3p^3q) 3/4 into radical form. I believe you meant What is the expression in radical form? (3p^3q) ^(3/4). In other words, the exponent of (3p^3q) is (3/4).

OpenStudy (mathmale):

What is the base here? The exponent? The power?

OpenStudy (mathmale):

\[In ~a ^{\frac{ b }{ c}}\] a is the base, b is the power and c is the index of the root.

OpenStudy (mouth4war):

index of the root?

OpenStudy (mathmale):

Can you now name a, b and c in \[(3p^3q) ^(3/4)?\]

OpenStudy (mathmale):

"index of the root:" examples:\[\sqrt{x}=\sqrt[2]{x}\]

OpenStudy (mathmale):

is the square root. The "index of the root" is 2 here.

OpenStudy (mathmale):

\[\sqrt[3]{x}=x ^{\frac{ 1 }{ 3 }}\]

OpenStudy (mathmale):

is the "cube root;" the "index of the root" is 3. And so on.

OpenStudy (mathmale):

Can you now name a, b and c in Can you now name a, b and c in (3p3q)^(3/4)?